y′ =y/x(x-x^3) ; y = −2 when x = 2




1
Expert's answer
2021-08-30T19:11:46-0400

"\displaystyle y' = \frac{y}{x(x-x^3)}
\\= \frac{y'}{y}= \frac{1}{x(x-x^3)}
\\\text{Next, we integrate the expression above, to do this we resolve $\frac{1}{x(x-x^3)}$ using partial }
\\\text{fractions}
\\\implies \frac{1}{x(x-x^3)}=\frac{1}{2(1-x)}+\frac{1}{2(1+x)}+\frac{1}{x^2}
\\\implies iny =\frac{1}{2}\int \frac{1}{1-x} + \frac{1}{2}\int \frac{1}{1+x} + \int \frac{1}{x^2}
\\= iny = -\frac{1}{2}in(1-x)+\frac{1}{2}in(1+x)-x^{-1}+c
\\\text{Taking the exponent of the expression above}
\\y = (\frac{1+x}{1-x})^\frac{1}{2}-e^{x^{-1}}+A
\\\text{Substituting y = -2, x= 2, we have that}
\\ A=-0.351-1.73i
\\\implies y = (\frac{1+x}{1-x})^\frac{1}{2}-e^{x^{-1}}-0.351-i.73i"